Spalding Pointers Corporation expects to begin operations on January 1, 2019, it will operate as a specialty sales company thDetermine the amount of sales revenue Spalding will report on the first 2019 quarterly pro forma income statement. Sales revePrepare a cash receipts schedule for the first quarter of 2019. January February March Schedule of Cash Receipts Receipts froDetermine the amount of accounts receivable as of March 31, 2019. Accounts receivable
Answer #1

1) Sales budget

January February March
Sales revenue 120000 120000*1.05 = 126000 126000*1.05 = 132300

2) Total sales revenue = $378300

3) Schedule of cash receipt

Schedule of cash receipt January February march
Receipt from January Sales 84000 24000 12000
Receipt from February Sales 88200 25200
Receipt from March Sales 92610
Total 84000 112200 129810

4) Account receivable = (126000*10%+132300*30%) = $52290